Tom has a 0.5 chance of winning the Booker Prize, whereas John has a 0.4 chance of winning the Booker Prize. Only one Booker prize is up for grabs. How likely is it that at least one of them will win the prize?


1
Expert's answer
2022-03-04T04:47:44-0500

Let A - "at least one of them will win the prize"

"H_1-" "Tom won"

"H_2-" "John won"

Then, "P(A)=P(H_1)*P(\neg H_2)+P(\neg H_1)*P(H_2)+P(H_1\cap H_2)=0.5*0.6+0.5*0.4+0=0.5"